OpenStudy (anonymous):

what is the simplified form of 8 times the square root of -27?

OpenStudy (anonymous):

So we have: \[ 8\sqrt{-27}=8\sqrt{-(9)(3)}=8\sqrt{-(3^2)(3)}=8(3)\sqrt{-3}=24\sqrt{-3} \]We know that: \[ i^2=-1 \]So: \[ 24\sqrt{-3}=24i\sqrt{3} \]Which is your final answer.